One number is 8 times a first number. A third number is 100 more than the first number. If the sum of the three numbers is 840, find the numbers
Answer by fractalier(6550)

Call the first number, x.
The second would be 8x.
The third would be x + 100.
Adding them we get
x + 8x + x+100 = 840
10x + 100 = 840
10x = 740
x = 74
8x = 592
x+100 = 174